# JEL Classifications: C61; D15; D52; E21; G51

### v1

@@ Abstract.
Empirical evidence shows that wealthy households have substantially higher saving rates and markedly lower marginal propensity to consume (MPC) than other groups. Existing theory can account for this pattern only under restrictive assumptions on returns, discounting, and preferences. This paper develops a general theory of optimal savings with preference shocks, allowing risk aversion to vary across states and over time. We show that incorporating such heterogeneity in risk attitudes fundamentally alters the asymptotic dynamics of consumption and saving. In particular, we provide an analytical characterization of the asymptotic MPCs and show that zero asymptotic MPCs, corresponding to a 100% asymptotic saving rate, arise under markedly weaker conditions than in existing theory. Strikingly, such outcomes occur whenever there is a positive probability that agents become less risk averse in the future. As a result, the vanishing MPC emerges as a generic feature rather than a knife-edge result of the optimal savings model, offering a more theoretically robust and empirically consistent account of the saving behavior of wealthy households.
@@
A Theory of Saving under Risk Preference Dynamics [^1] Qingyin Ma gratefully acknowledges financial support from the National Natural Science Foundation of China (Grant No. 72573111).
@@ 1. Introduction
In recent decades, a large body of empirical evidence has shown that wealthy households save a substantially larger fraction of their wealth and exhibit markedly lower marginal propensity to consume (MPC) than other groups. For instance, Dynan et al. (2004) document a strong positive relationship between lifetime income and saving rates. Using Norwegian administrative data, Fagereng et al. (2025) find that households in the top 1% of the wealth distribution report saving rates far exceeding those of the median. Because household saving behavior directly drives wealth accumulation, and wealthy households hold a disproportionate share of total assets (Saez and Zucman, 2016), their saving patterns have first-order implications for the long-run distribution of wealth and the mechanisms that sustain inequality (Carroll et al., 2017). From a macroeconomic perspective, even small variations in the MPC or risk-taking behavior of the wealthy can have outsized effects on aggregate saving rates, capital formation, asset prices, the transmission of fiscal and financial shocks, the equilibrium interest rate, and the design of macroprudential policy (Ampudia et al., 2024; Carroll et al., 2017; Cho, 2023; Crawley and Kuchler, 2023; Jappelli and Pistaferri, 2014; Mian et al., 2021; Patterson, 2023).
@@
Despite these empirical regularities, existing theoretical work remains limited in explaining the persistent high saving rates and low MPCs of wealthy households. [^2] As is standard in the literature, we define the saving rate as the change in net worth relative to total income, excluding capital losses: $s_{t+1}=\frac{w_{t+1}-w_{t}}{\max\{(R_{t+1}-1)(w_{t}-c_{t}),0\}+Y_{t+1}},$ where $s_{t}$, $w_{t}$, $c_{t}$, $Y_{t}$, and $R_{t}$ denote the saving rate, wealth, consumption, nonfinancial income, and the gross rate of return on wealth, respectively. Moreover, the asymptotic saving rate refers to the saving rate in the limit as wealth $w_{t}\to\infty$. The canonical incomplete market model, as emphasized by Ma and Toda (2021), predicts that in a stationary equilibrium with constant relative risk aversion, discount factor, and return on wealth, the asymptotic saving rate is negative, contradicting the empirical evidence that the wealthy continue to accumulate assets indefinitely. Subsequent studies have introduced return risk and stochastic discounting to reconcile theory with data, yet these extensions largely preserve the assumption of constant risk aversion (Benhabib et al., 2015; Ma et al., 2020; Ma and Toda, 2021). Even in richer stochastic environments, as Ma and Toda (2021) show, matching the observed low MPCs and high saving rates among wealthy households requires quantitatively stringent conditions, and the resulting implications remain fragile.
@@
To begin, we study a generic optimal savings problem in which agents face preference shocks, and the rate of return, discount factor, and nonfinancial income are all allowed to be state-dependent, time-varying, and mutually dependent. In this setting, the state space for wealth can be unbounded, and the reward function can be unbounded above and/or below, making it challenging to apply standard dynamic programming techniques in the usual space of value functions. [^3] A further complication is that characterizing the value function as a fixed point of the Bellman operator provides limited insight into the properties of the optimal consumption policy. To overcome these challenges, we follow the approach of Li and Stachurski (2014) and Ma et al. (2020), introducing a supremum distance between marginal utilities, and constructing an operator directly from the Euler equation. We then show that the optimal policy exists as the unique fixed point of this operator in a suitable candidate space. Our key assumption admits a natural interpretation in terms of the discounted average gross payoff on assets, which can be easily verified in practice.
@@
Our theory therefore provides a more theoretically robust and empirically consistent account of the saving behavior of wealthy households, highlighting how risk preference heterogeneity offers a natural mechanism for the persistence of wealth accumulation and inequality.
@@
Our study also builds on the theoretical literature on the optimal savings problem, a central building block of incomplete market models in modern macroeconomics (Chamberlain and Wilson, 2000; Rabault, 2002; Kuhn, 2013; Li and Stachurski, 2014; Benhabib et al., 2015; Ma et al., 2020). As noted earlier, we extend the techniques of Li and Stachurski (2014) and Ma et al. (2020) to solve a generic optimal savings problem with preference shocks. While this literature focuses primarily on existence, uniqueness, and computability, our analysis centers on the asymptotic properties of optimal consumption, which has received little attention and has no direct counterpart in these studies.
Finally, our model is motivated by the empirical evidence that risk aversion is time-varying and state-dependent, evolving with both idiosyncratic and aggregate shocks. Improvements in financial circumstances and past success are associated with lower risk aversion (Andersen et al., 2008; Kettlewell, 2019; Yesuf and Bluffstone, 2009). Adverse events such as health shocks (Deckere and Schmitz, 2016; Finkelstein et al., 2013), exposure to violence (Jakiela and Ozier, 2019), or accidents (Shum and Xin, 2022) often lead to substantial and persistent increases in risk aversion. Macroeconomic conditions also play a crucial role: risk aversion tends to decline during expansions (Sahm, 2012; Malmendier and Nagel, 2011) and rise sharply following recessions, such as the global financial crisis (Guiso et al., 2018).
The rest of the paper is organized as follows. Section 2 formulates a generic optimal savings problem with preference shocks, and establishes existence, uniqueness, and key structural properties of the optimal policy. Section 3 systematically studies its asymptotic properties and discusses the resulting implications for saving and wealth accumulation. Section 4 concludes. Proofs are relegated to the Appendix.
@@ 3.2. The Asymptotic MPCs
Interpretation Less risk averse with positive probability Same risk aversion with probability 1 Strictly more risk averse with probability 1 Weakly more risk averse with positive probability
@@
$\bar{c}(z)=\begin{cases}1-(\mathbbm{E}\beta R^{1-\gamma})^{1/\gamma},&\text{if }\mathbbm{E}\beta R^{1-\gamma}<1,\\ 0,&\text{otherwise.}\end{cases}$
@@ 4. Conclusion
We developed a generalized optimal savings model with preference shocks, allowing for time-varying and state-dependent risk aversion. This extension fundamentally reshapes the asymptotic behavior of consumption and saving. Our analysis shows that zero asymptotic MPCs arise naturally whenever agents face a positive probability of becoming less risk averse, even in the absence of return risk or stochastic discounting. This mechanism provides a theoretically robust and empirically consistent explanation for the persistent accumulation of wealth and the high saving rates observed among the rich. By linking saving behavior to the dynamics of risk preferences, our framework highlights a new channel through which heterogeneity in attitudes toward risk could influence aggregate saving and account for wealth inequality and asset market outcomes. We leave these topics for future research.

### v2

@@ Abstract.
Empirical evidence shows that wealthy households have substantially higher saving rates and markedly lower marginal propensity to consume (MPC) than other groups. Existing theory cannot account for this pattern unless under restrictive assumptions on returns, discounting, and preferences. This paper develops a general theory of optimal savings with preference shocks, allowing risk aversion to vary across states and over time. We show that incorporating such heterogeneity in risk attitudes fundamentally alters the asymptotic dynamics of consumption and saving. In particular, we provide an analytical characterization of the asymptotic MPCs and show that zero asymptotic MPCs, corresponding to a 100% asymptotic saving rate, arise under markedly weaker conditions than in existing theory. Strikingly, such outcomes occur whenever there is a positive probability that agents become less risk averse in the future. As a result, the vanishing MPC emerges as a generic feature rather than a knife-edge result of the optimal savings model, offering a more theoretically robust and empirically consistent account of the saving behavior of wealthy households.
@@
A Theory of Saving under Risk Preference Dynamics [^1] We thank Shenghao Zhu, Nisha Peng, John Stachurski, and seminar audiences at CUEB and CUFE for helpful comments and suggestions. Qingyin Ma gratefully acknowledges financial support from the National Natural Science Foundation of China (Grant No. 72573111).
@@ 1. Introduction
In recent decades, a large body of empirical evidence has shown that wealthy households save a substantially larger fraction of their wealth and exhibit markedly lower marginal propensity to consume (MPC) than other groups. For instance, Dynan et al. (2004) document a strong positive relationship between lifetime income and saving rates. Using Norwegian administrative data, Fagereng et al. (2025) find that households in the top 1% of the wealth distribution report saving rates far exceeding those of the median.
Because household saving behavior directly drives wealth accumulation, and wealthy households hold a disproportionate share of total assets, [^2] Saez and Zucman (2016) document that the top 0.1% wealth share in the United States increased from 7% in 1979 to 22% in 2012. Similarly, Piketty et al. (2019) estimate that the top 10% wealth share in China rose from 40% in 1995 to 67% in 2015. their saving patterns have first-order implications for the long-run distribution of wealth and the mechanisms that sustain inequality (Carroll et al., 2017). From a macroeconomic perspective, even small variations in the MPC or risk-taking behavior of the wealthy can have outsized effects on aggregate saving rates, capital formation, asset prices, the transmission of fiscal and financial shocks, the equilibrium interest rate, and the design of macroprudential policy (Ampudia et al., 2024; Carroll et al., 2017; Cho, 2023; Crawley and Kuchler, 2023; Jappelli and Pistaferri, 2014; Mian et al., 2021; Patterson, 2023).
@@
Despite these empirical regularities, existing studies remain limited in explaining the persistent high saving rates and low MPCs of wealthy households. [^3] As is standard in the literature, we define the saving rate as the change in net worth relative to total income, excluding capital losses: $s_{t+1}=\frac{w_{t+1}-w_{t}}{\max\{(R_{t+1}-1)(w_{t}-c_{t}),0\}+Y_{t+1}},$ where $s_{t}$, $w_{t}$, $c_{t}$, $Y_{t}$, and $R_{t}$ denote the saving rate, wealth, consumption, nonfinancial income, and the gross rate of return on wealth, respectively. Moreover, the asymptotic saving rate refers to the saving rate in the limit as wealth $w_{t}\to\infty$. The canonical incomplete market model, as emphasized by Ma and Toda (2021), predicts that in a stationary equilibrium with constant relative risk aversion, discount factor, and return on wealth, the asymptotic saving rate is negative, contradicting the empirical evidence that the wealthy continue to accumulate assets indefinitely. Subsequent studies have introduced return risk and stochastic discounting to reconcile theory with data, yet these extensions largely preserve the assumption of constant risk aversion (Benhabib et al., 2015; Ma et al., 2020; Ma and Toda, 2021). Even in richer stochastic environments, as Ma and Toda (2021) show, matching the observed low MPCs and high saving rates among wealthy households requires quantitatively stringent conditions, and the resulting implications remain fragile.
@@
To begin, we study a generic optimal savings problem, in which agents face preference shocks, and the rate of return, discount factor, and nonfinancial income are all allowed to be state-dependent, time-varying, and mutually dependent. We establish the existence and uniqueness of the optimal consumption policy in a suitable candidate space. Our key assumption admits a transparent interpretation in terms of the discounted average gross payoff on assets and is straightforward to verify in practice.
@@
We then complement our theoretical analysis with a quantitative investigation. First, we provide numerical illustrations that render the asymptotic results transparent and economically intuitive. Second, following Koop et al. (1996), we compute state-dependent impulse response functions to uncover the economic mechanisms through which preference shocks reshape consumption and saving dynamics.
Our quantitative analysis confirms that shocks to risk aversion generate economically significant and persistent responses in household consumption and saving. Importantly, both the magnitude and persistence of these responses depend on the household’s current state and on the persistence of preference shocks. In particular, the effects vary systematically with wealth and the prevailing level of risk aversion, revealing pronounced nonlinearities in the transmission of preference dynamics across the wealth distribution.
Taken together, our framework provides a theoretically robust and empirically consistent account of the saving behavior of wealthy households, highlighting how risk preference heterogeneity offers a natural mechanism for the persistence of wealth accumulation. While we do not explicitly model wealth inequality, our analysis points to a novel mechanism that may be relevant for understanding wealth accumulation in richer environments.
@@
Our study also builds on the theoretical literature on the optimal savings problem, a central building block of incomplete market models in modern macroeconomics (Chamberlain and Wilson, 2000; Rabault, 2002; Kuhn, 2013; Li and Stachurski, 2014; Benhabib et al., 2015; Ma et al., 2020). We extend the techniques of Li and Stachurski (2014) and Ma et al. (2020) to solve a generic optimal savings problem with preference shocks. While this literature focuses primarily on existence, uniqueness, and computability, our analysis centers on the asymptotic properties of optimal consumption, which has received little attention and has no direct counterpart in these studies.
The rest of the paper is organized as follows. Section 2 formulates a generic optimal savings problem with preference shocks, and establishes existence, uniqueness, and key structural properties of the optimal policy. Section 3 systematically studies its asymptotic properties and discusses the resulting implications for saving and wealth accumulation. Section 4 presents a quantitative analysis that provides further intuition for our theoretical results. Section 5 concludes. Proofs are relegated to the Appendix.
@@ 3.2. The Asymptotic MPCs
Interpretation Less risk averse with positive probability Generalized knife-edge condition Same risk aversion with probability 1 Strictly more risk averse with probability 1 Weakly more risk averse with positive probability
@@
$\bar{c}(z)=\begin{cases}0,&\text{if }\mathbbm{E}\beta R^{1-\gamma}\geq 1\\ 1-(\mathbbm{E}\beta R^{1-\gamma})^{1/\gamma},&\text{otherwise.}\end{cases}$
@@ 3.3. The Asymptotic Saving Rate
4. Quantitative Analysis
The preceding theoretical analysis shows that allowing for dynamically evolving risk aversion fundamentally alters the consumption–saving dynamics of the model. In this section, we complement the theory with a quantitative analysis that serves two purposes. First, using a parsimonious quantitative specification, we provide numerical illustrations that render the main theoretical results transparent and economically intuitive. Second, within a more general quantitative environment, we compute generalized impulse response functions to elucidate the economic mechanisms through which preference shocks generate qualitative changes in model dynamics.
4.1. Model Specification
For our quantitative analysis, we consider an optimal savings problem with shocks to risk aversion. The discount factor and the gross rate of return on wealth are held constant over time, with $\beta_{t}\equiv\beta$ and $R_{t}\equiv R$. Risk aversion follows a discretized autoregressive process of order one:
$\gamma_{t+1}=\mu_{\gamma}(1-\rho_{\gamma})+\rho_{\gamma}\gamma_{t}+\sigma_{\gamma}\sqrt{1-\rho_{\gamma}^{2}}\varepsilon_{t+1}^{\gamma},\quad(\varepsilon_{t}^{\gamma})\mathrel{{\mathop{\sim}\limits^{\textrm{ {\sc iid }}}}}N(0,1).$
This specification is a special case of our theoretical framework, with the preference state given by $\bar{Z}_{t}=\gamma_{t}$. For the nonfinancial income process, we allow for alternative specifications depending on the quantitative experiment under consideration. The model is solved using the endogenous grid method of Carroll (2006). Below we describe the key elements of the quantitative setup. Detailed implementation and algorithmic steps are provided in Appendix E.
In our quantitative experiments, we work at an annual frequency and set $\beta=0.95$ and $R=1.02$. In our baseline setting, the risk aversion process is parameterized by $\rho_{\gamma}=0.6$, $\mu_{\gamma}=4$, and $\sigma_{\gamma}=1$, and discretized into a finite-state Markov chain using the method of Tauchen (1986). The number of discrete states is allowed to vary across experiments. In particular, for the impulse response analysis in Section 4.3, we employ a relatively fine discretization to improve approximation accuracy. Across all specifications, the minimum, median (and mean), and maximum values of $\gamma_{t}$ are $1$, 44, and $7$, respectively, which lie well within the range commonly used in empirical research.
4.2. The Asymptotic Saving Rates
To better illustrate the implications of our theory, we begin by comparing our model with the standard optimal savings problem without preference shocks. To this end, we consider simply $Y_{t}\equiv 1$ and discretize $\{\gamma_{t}\}$ into a three-state Markov chain. The resulting state space and transition matrix are
$\gamma_{t}\in\{1,4,7\}\quad\text{and}\quad\bar{P}=\begin{bmatrix}0.6416&0.3538&0.0001\\ 0.0303&0.9392&0.0305\\ 0.0001&0.3538&0.6461\end{bmatrix}$.
The stationary mean of this discrete process coincides with the median state $\gamma_{t}=4$. At this state, the probability of remaining at $\gamma_{t}=4$ next period is 93.92%, while the probability of transitioning to the lower risk aversion state $\gamma_{t}=1$ is only 3.03%.
Figure 1. The optimal consumption function
To ensure a fair comparison, we contrast the optimal consumption function when $\gamma_{t}$ is fixed at its stationary mean $\mathbbm{E}\gamma_{t}=4$ in our stochastic risk aversion model with that of the standard optimal savings problem featuring CRRA utility with a constant risk aversion coefficient $\gamma=4$. Figure 1 plots the results. Despite the fact that current risk aversion is identical across the two models and that there is a high probability of remaining at the same level of risk aversion next period, consumption is substantially lower under stochastic risk aversion (red dotted line) than under constant risk aversion (blue solid line), reflecting the precautionary behavior induced by the possibility of future changes in risk aversion.
Figure 2. The consumption and saving rates
Figure 2 plots the consumption rate, defined as $c^{*}(w,z)/w$, and the saving rate as a function of wealth. The left panel shows that, in the standard optimal savings model, the consumption rate (blue solid line) converges to the theoretical asymptotic MPC, which is strictly positive and equal to $1-(\beta R^{1-\gamma})^{1/\gamma}=0.0273$ as derived in Example 3.5. By contrast, in the stochastic risk aversion model, the consumption rate exhibits a pronounced downward trend and continues to decline even at very high levels of wealth. This pattern is fully consistent with Theorem 3.1, which establishes that the asymptotic MPC is equal to zero.
The right panel plots the corresponding saving rates. In the standard optimal savings model, saving rates are negative and converge to their theoretical asymptotic value of $40.31\%$, as given by (17), in stark contradiction with the empirical evidence on wealthy households. In contrast, under stochastic risk aversion, the saving rate converges to its theoretical limit of $100\%$. Taken together, these results illustrate that even small probabilities of shifts in risk aversion can have sizable effects on the saving behavior, suggesting that stochastic risk aversion offers a promising mechanism for generating persistently high saving rates among wealthy households, in line with the patterns observed in the data.
4.3. Impulse Response Analysis
In this section, we systematically examine how changes in risk aversion influence consumption and saving behavior using impulse response functions (IRFs). To capture the model’s nonlinear dynamics, we construct IRFs as state-dependent random variables, following Koop et al. (1996). These IRFs allow us to trace how shocks to risk aversion propagate over time and affect consumption, saving rates, and consumption volatility across different states. [^11] Consumption volatility in period-$t$ is defined as the conditional variance: $\operatorname{Var}_{t-1}[c^{*}(w_{t},Z_{t})]:=\mathbbm{E}_{t-1}[c^{*}(w_{t},Z_{t})]^{2}-[\mathbbm{E}_{t-1}c^{*}(w_{t},Z_{t})]^{2}.$
We follow the standard earnings specification in the literature and assume that labor income evolves according to
$\log Y_{t}=\eta_{t}+\nu_{t},\quad(\nu_{t})\sim N(0,\sigma_{v}^{2}),$
$\eta_{t+1}=\rho_{\eta}\eta_{t}+\varepsilon_{t}^{\eta},\quad(\varepsilon_{t}^{\eta})\sim N(0,\sigma_{\eta}^{2}).$
Consistent with the empirical estimates reported in Lander (2024), we set $\rho_{\eta}=0.9735$, $\sigma_{\eta}=0.1272$, and $\sigma_{\nu}=0.2102$, based on PSID data over 1970-2012.
To ensure sufficient numerical accuracy, we discretize $\{\gamma_{t}\}$ into a 121-state Markov chain following Tauchen (1986). The persistent income component $\{\eta_{t}\}$ is discretized into 5 states, while expectations over the transitory shock $\{\nu_{t}\}$ are computed using a 7-point Gauss–Hermite quadrature. Sensitivity analysis confirms that the results are robust to these parametric choices.
Figure 3 reports impulse response functions to a one-unit increase in risk aversion at the beginning of period $t$. As noted earlier, the IRFs are computed as state-dependent responses, conditional on different initial levels of wealth, risk aversion, and the persistent income component $(w_{t-1},\gamma_{t-1},\eta_{t-1})$. All experiments are conducted conditional on $\eta_{t-1}=0$, its stationary mean.
The top row of Figure 3 illustrates heterogeneity with respect to initial wealth, holding risk aversion fixed at its stationary mean $\gamma_{t-1}=4$. The left panel shows that an increase in risk aversion induces an immediate and persistent decline in consumption, with the magnitude of the response rising sharply in wealth. When wealth rises from $w_{t-1}=5$ to $10$ and $50$, contemporaneous consumption falls by 4.67%, 7.89%, and 20.23%, respectively, with effects lasting roughly 2–6 years. This nonlinearity reflects the fact that low-wealth households are constrained in their ability to adjust savings, so consumption responds weakly to preference shocks, whereas wealthy households can reallocate intertemporal resources more flexibly.
The top-middle panel shows that this adjustment operates primarily through saving. A one-unit increase in risk aversion raises the saving rate by 5.81%, 10.15%, and 28.05% at $w_{t-1}=5,10,$ and $50$, respectively, with effects persisting for 4–8 years. Thus, as wealth rises, the response shifts from limited consumption adjustment toward substantial increases in saving. This pattern aligns with our theoretical mechanism: at high wealth levels, precautionary motives dominate, prompting agents to accumulate assets aggressively even in response to transitory risk aversion shocks. The top-right panel further shows that consumption volatility declines following a positive risk aversion shock, particularly for wealthier households, reflecting stronger consumption smoothing driven by heightened precautionary behavior.
The middle row of Figure 3 illustrates heterogeneity with respect to initial risk aversion, holding initial wealth fixed at $w_{t-1}=10$. When initial risk aversion is high ($\gamma_{t-1}=6$), the same shock generates relatively modest responses: consumption falls by 4.84% and the saving rate rises by 5.53%. By contrast, when agents start from a low risk aversion state ($\gamma_{t-1}=1$), consumption drops sharply by 19.26% and saving increases by 32.21%. This asymmetry reflects state-dependent amplification: a positive shock represents a larger proportional change in effective prudence when baseline risk aversion is lower, triggering a stronger re-optimization of intertemporal choices. Consistent with this mechanism, the decline in consumption volatility is also largest at low initial risk aversion.
Figure 3. Impulse responses to a positive one-unit risk aversion shock
The bottom row of Figure 3 examines the role of the persistence of risk aversion shocks, governed by $\rho_{\gamma}$, holding initial wealth and risk aversion fixed at $(w_{t-1},\gamma_{t-1})=(10,4)$. Higher persistence markedly amplifies and prolongs the responses of consumption and saving. As $\rho_{\gamma}$ increases, households cut consumption more sharply, raise saving more persistently, and experience larger declines in consumption volatility. Intuitively, greater persistence raises the expected duration of elevated risk aversion, strengthening precautionary motives and leading agents to treat preference shocks as quasi-permanent. This persistence channel thus reinforces the state-dependent and nonlinear dynamics emphasized by our theory.
Taken together, these impulse responses deliver three key insights. First, shocks to risk aversion generate economically large and persistent movements in consumption and saving. Second, these effects are highly state dependent, varying systematically with households’ wealth, the current level of risk aversion, and the persistence of preference shocks. Third, the IRFs provide a quantitative illustration of the core theoretical mechanism: stochastic risk preferences alone can fundamentally reshape saving behavior, even in environments with constant returns and discounting, accounting for the high and persistent saving rates observed among wealthy households.
@@ 5. Conclusion
We developed a generalized optimal savings model with preference shocks, allowing for time-varying and state-dependent risk aversion. This extension fundamentally reshapes the asymptotic behavior of consumption and saving. Our analysis shows that zero asymptotic MPCs arise naturally whenever agents face a positive probability of becoming less risk averse, even in the absence of return risk or stochastic discounting.
Quantitative illustrations further confirm that this mechanism can generate large and persistent variations in consumption and saving across states, consistent with observed wealth accumulation patterns. Taken together, our framework provides a theoretically robust and empirically consistent explanation for the persistent accumulation of wealth and the high saving rates observed among the rich.
By linking saving behavior to the dynamics of risk preferences, our model highlights a potential channel through which heterogeneous risk attitudes may influence aggregate saving and offers insights relevant for understanding wealth inequality and asset market outcomes. A detailed exploration of these implications is left for future research.